# Novel inner shadows of the Kerr black hole illuminated by a tilted thin accretion disk

### v1

@@ Abstract
Inner shadow; Kerr black hole; Tilted accretion disk; Ray-tracing
I Introduction
@@
Since the advent of general relativity, black holes have remained a subject of profound fascination within the astronomical community. In recent years, the study of black holes has evolved from theoretical speculation to direct observation. Notably, the Event Horizon Telescope (EHT) collaboration, utilizing a Very Long Baseline Interferometry (VLBI) array, successfully captured and released the first images of the supermassive black holes at the centers of the M87 galaxy and the Milky Way Akiyama et al. (2019) ; Akiyama et al. (2022) . These landmark observations not only conclusively verified the existence of black holes but also opened a novel channel for their exploration. Subsequently, a substantial body of research on black hole images Narayan et al. (2019) ; Gralla et al. (2019) ; Bronzwaer and Falcke (2021) ; Chael et al. (2021) ; Lin et al. (2022) ; Vincent et al. (2022) ; Hou et al. (2022) ; Heydari-Fard et al. (2023a) ; Heydari-Fard et al. (2023b) ; Zhang et al. (2023) ; Hu et al. (2023) ; Hu et al. (2024) ; Wang et al. (2024) ; Li et al. (2024b) , black hole-jet system images Chatterjee et al. (2020) ; Kawashima et al. (2021) ; Lu et al. (2023) ; Papoutsis et al. (2023) ; Zhang et al. (2024a) , hot-spot images Tamm and Rosa (2024) ; Rosa et al. (2024) ; Huang et al. (2024) ; Chen et al. (2024) ; Zhou et al. (2024) ; Wu et al. (2024) , and polarized images Zhang et al. (2021) ; Gelles et al. (2021) ; Qin et al. (2022) ; Hu et al. (2022b) ; Zhang et al. (2024b) ; Shi et al. (2024) ; Guo et al. (2024) , within the frameworks of general relativity and modified gravity theories, has emerged, injecting unprecedented vitality into this field.
@@
Theoretically, a black hole can capture light rays emitted by surrounding sources, resulting in a distinct region of brightness depression on the observation plane from Earth. This region is referred to as the ‘shadow’, first introduced by Falcke et al. Falcke et al. (2000) , and its boundary is the 2D projection of the black hole’s unstable critical photon orbits. Notably, the black hole shadow is considered the ‘fingerprint’ of spacetime, unaffected by the accretion environment, making it a powerful tool for studying black hole physics and gravity theories. Consequently, many pioneering studies have focused on constraining system parameters through the shadow’s profile Hioki and Maeda (2009) ; Johannsen (2013) ; Wei et al. (2019) ; Gralla et al. (2020) ; Li et al. (2020) ; Farah et al. (2020) ; Hou et al. (2021) ; Afrin and Ghosh (2022) ; Cao et al. (2023) . However, accurately extracting the geometric information of the black hole shadow requires extremely high angular resolution, which is currently beyond reach with the technology of the EHT Gralla et al. (2019) ; Gralla (2021) . As a result, despite the groundbreaking achievements of the EHT, there remains ample room for discussion regarding modifications to gravity theories and the accretion environment.
@@
On the other hand, it is interesting to note that for some supermassive black holes at the centers of galaxies, such as Sagittarius A, the accreting material near the event horizon is optically thin at $230$ GHz Akiyama et al. (2022) . In other words, following the backward ray-tracing approach, the accretion flow within the innermost stable circular orbit (ISCO) can intercept light rays that would otherwise pass through the black hole’s event horizon, allowing the observer to discern a dim region smaller than the traditional shadow. Numerous studies have confirmed that such a dim region can appear when the inner edge of an equatorial, optically and geometrically thin accretion disk coincides with the event horizon, and its boundary is clearly separated from the higher-order bright rings Peng et al. (2021) ; Li and He (2021) ; Hu et al. (2022a) ; Zeng et al. (2022) ; Guo et al. (2022) ; Gao et al. (2023) ; Meng et al. (2023) ; Wang et al. (2023) ; Yang et al. (2023) ; Li et al. (2024a) ; Sui et al. (2024) . More importantly, Chael et al. identified a similar region in general relativistic magnetohydrodynamic (GRMHD) simulations in Kerr spacetime, which they named the ‘inner shadow’ to distinguish it from the traditional shadow Chael et al. (2021) . Furthermore, researchers have pointed out that the boundary of the inner shadow is a projection of the black hole’s event horizon, and its geometric informations have the potential to constrain system parameters such as the viewing angle and spin. Building on this, several authors have explored the geometric properties of the inner shadow of Kerr-Melvin black holes illuminated by an equatorial accretion disk across different parameter spaces, finding that the evolution of the inner shadow with varying parameters aligns with the changes in higher-order photon rings. This has led them to propose that the black hole’s inner shadow could serve as a tool for identifying black holes and studying the magnetic field around them Hou et al. (2022) .
@@
In astrophysics, the angular momentum of accreting matter is not always aligned with the angular momentum of the black hole, which leads to the formation of tilted accretion disk. There is substantial evidence showing that tilted accretion disks are commonly present in X-ray binary systems Maccarone (2002) ; Caproni et al. (2006) ; van den Eijnden et al. (2017) , and the tilted accretion disk-jet system of the supermassive black hole M87 has also been confirmed Cui et al. (2023) . Research on the structure evolution and energy spectrum of tilted accretion disks has garnered considerable attention in the scientific community Mckinney et al. (2013) ; Liska et al. (2019) ; Liska et al. (2021) ; Musoke et al. (2023) ; Liska et al. (2023) . Inspired by this, Hu and colleagues conducted a simulation of the images of spherically symmetric hairy black holes illuminated by a geometrically thin, tilted accretion disk within the framework of Horndeski theory Hu et al. (2024) . They found that the inner shadow of the black hole tends to be obscured by the tilted accretion disk, and it also rotates along with the precession of the disk. Therefore, they argued that, given the difficulties in constraining the accretion environment, using the inner shadow to test gravity theories requires particular caution. However, the scientific community has reached a consensus on the existence of rotation in astrophysical black holes, indicating that Hu et al. did not tell the whole story. In the scenario of a rotating black hole, does a tilted accretion disk obscure the inner shadow? What insights can the inner shadow provide for black hole physics and high-energy physics? These questions will be properly addressed in this paper.
@@
The remainder of this paper is organized as follows. In Sec. II, we briefly review the Kerr black hole and the canonical equation of photon motion, which form the foundation of the ray-tracing algorithm. Then, in Sec. III, we provide a detailed introduction to the backward ray-tracing method and the model of the tilted accretion disk. Before the conclusions and discussion in Sec. V, we present in Sec. IV the novel inner shadows cast by the Kerr black hole illuminated by a tilted thin accretion disk across different parameter spaces, and carefully elaborate on their astrophysical implications. Throughout this paper, we adopt geometric units, where the black hole mass $M$, the speed of light $c$, and the gravitational constant $G$ are set to unity.
II Kerr metric and null-like geodesics
@@
Astrophysical black holes are usually described by the Kerr metric, which can be expressed in the Boyer-Lindquist coordinates $x=(t,r,\theta,\varphi)$ with the spacelike signature (,+,+,+)(-,+,+,+)( - , + , + , + ) as Papoutsis et al. (2023) ; Hioki and Maeda (2009) ; Cunha et al. (2016) ; Pu et al. (2016) ; Cadavid et al. (2022)
@@
$\textrm{d}s^{2}$ $=$ $-\left(1-\frac{2r}{\Sigma}\right)\textrm{d}t^{2}-\frac{4ar\sin^{2}\theta}{\Sigma}\textrm{d}t\textrm{d}\varphi+\frac{\Sigma}{\Delta}\textrm{d}r^{2}+\Sigma\textrm{d}\theta^{2}$ (1)
@@
$\Sigma=r^{2}+a^{2}\cos^{2}\theta,$ (2)
$\Delta=r^{2}-2r+a^{2}.$ (3)
@@
$r_{\textrm{e}}=1+\sqrt{1-a^{2}}.$ (4)
It is important to emphasize that the size of the inner shadow of a black hole is associated with the event horizon Chael et al. (2021) . Consequently, it is anticipated that the inner shadow of a Kerr black hole shrinks as the spin parameter increases. Given that there is an upper limit to the spin parameter for a Kerr black hole solution, the size of the inner shadow theoretically has a minimum value. We will delve into this topic in detail in Sec. IV.
@@
$\mathscr{H}=p_{\mu}\dot{x^{\mu}}-\mathscr{L}=\frac{1}{2}g^{\mu\nu}p_{\mu}p_{\nu},$ (6)
@@
$p_{t}=-\left(1-\frac{2r}{\Sigma}\right)\dot{t}-\frac{2ar\sin^{2}\theta}{\Sigma}\dot{\varphi},$ (7)
$p_{r}=\frac{\Sigma}{\Delta}\dot{r},$ (8)
$p_{\theta}=\Sigma\dot{\theta},$ (9)
$p_{\varphi}=\sin^{2}\theta\left(r^{2}+a^{2}+\frac{2a^{2}r\sin^{2}\theta}{\Sigma}\right)\dot{\varphi}-\frac{2ar\sin^{2}\theta}{\Sigma}\dot{t}.$ (10)
It is necessary to point out that since the Kerr spacetime is of Petrov type D, the metric is independent of the coordinates $t$ and $\varphi$. Consequently, $p_{t}$ and $p_{\varphi}$ are conserved quantities in photon motion, corresponding to the specific energy $E=-p_{t}$ and specific angular momentum $L=p_{\varphi}$, respectively. Hence, we obtain the specific expression of (6) as
$\mathscr{H}=\frac{1}{2}\left[-\frac{\left(\Delta+2r\right)^{2}-\Delta a^{2}\sin^{2}\theta}{\Delta\Sigma}(-E)^{2}+\frac{\Delta}{\Sigma}p_{r}^{2}+\frac{p_{\theta}^{2}}{\Sigma}+\frac{\Sigma-2r}{\Delta\Sigma\sin^{2}\theta}L^{2}-\frac{4ar}{\Delta\Sigma}(-E)L\right].$ (11)
@@
$\dot{x^{\mu}}=\frac{\partial\mathscr{H}}{\partial p_{\mu}},\quad\dot{p_{\mu}}=-\frac{\partial\mathscr{H}}{\partial x^{\mu}},$ (12)
@@ III Ray-tracing method
Fig. 1: (colour online) Diagram of the ray-tracing method. The local reference frames of the black hole and the observer are denoted by $(x^{\prime},y^{\prime},z^{\prime})$ and $(x,y,z)$, respectively, with the observer’s screen located in the $\overline{xoy}$ plane. $\Theta$ and $\Phi$ represent the observer’s inclination angle and azimuthal angle, respectively. The disk, idealized as a tilted thin accretion disk, is composed of four distinct color-coded quadrants, with its inner boundary connected to the event horizon of the black hole. The tilt angle $\sigma$ is defined as the angle between the black hole’s equatorial plane $\overline{x^{\prime}o^{\prime}y^{\prime}}$ and the tilted accretion disk. Following the positive direction of $\Phi$, the quadrants of the disk are colored red, green, blue, and yellow, respectively. By employing numerical integration algorithms, the trajectories of light rays originating from each pixel $(x,y,0)$ on the observer’s screen can be tracked, with those terminating at the event horizon forming the inner shadow.
As shown in Fig. 1, the local frames of the observer and the Kerr black hole are represented by quasi-Cartesian coordinate systems $(x,y,z)$ and $(x^{\prime},y^{\prime},z^{\prime})$, respectively. Here, the $z$-axis coincides with $\overline{oo^{\prime}}$ and points towards the point $o^{\prime}$, with an angle $\Theta$ relative to the black hole’s rotation axis ($z^{\prime}$-axis), which defined as the viewing angle. The observer’s screen, depicted as a purple block, lies in $\overline{xoy}$ plane, thus $(x,y,0)$ naturally denotes the initial position of the photon. The solid disk inclined at an angle $\sigma$ to the black hole’s equatorial plane represents a tilted, geometrically thin accretion disk. It is divided into four regions of different colors by quadrants to vividly illustrate the black hole’s inner shadow and the spacetime distortion. It is important to note that for the equatorial accretion disk scenario $(\sigma=0^{\circ})$, the black hole image is independent of the observer’s azimuthal angle. However, this property is expected to not hold when a non-zero $\sigma$ is introduced. Therefore, an additional parameter, the observation azimuth $\Phi$, is included in our model.
Next, we derive the initial position of light rays in the black hole’s local coordinate system. From Fig. 1, it can be observed that under the condition of asymptotically flatness of spacetime, the observer’s local coordinate system $(x,y,z)$ can be aligned with the black hole’s local coordinate system $(x^{\prime},y^{\prime},z^{\prime})$ through rotation and translation. Thus, the light ray emitted from a pixel $(x,y,z)$ on the observer’s plane can be expressed in the black hole’s local coordinate system as Lin et al. (2022) ; Pu et al. (2016) ; Younsi et al. (2016)
$x^{\prime}$ $=$ $\mathscr{T}\cos\Phi-x\sin\Phi,$ (13)
$y^{\prime}$ $=$ $\mathscr{T}\sin\Phi+x\cos\Phi,$ (14)
$z^{\prime}$ $=$ $\left(r_{\textrm{obs}}-z\right)\cos\Theta+y\sin\Theta,$ (15)
@@
$\mathscr{T}=\left(\sqrt{r_{\textrm{obs}}^{2}+a^{2}}-z\right)\sin\Theta-y\cos\Theta.$ (16)
@@
$r$ $=$ $\sqrt{\frac{\mathscr{D}+\sqrt{\mathscr{D}^{2}+4a^{2}z^{\prime 2}}}{2}},$ (17)
$\theta$ $=$ $\arccos\left(\frac{z^{\prime}}{r}\right),$ (18)
$\varphi$ $=$ $\textrm{atan}2\left(y^{\prime},x^{\prime}\right),$ (19)
@@
$\mathscr{D}=x^{\prime 2}+y^{\prime 2}+z^{\prime 2}-a^{2}.$ (20)
To integrate the geodesic equations, in addition to determining the initial position of the photon, it is also necessary to specify its covariant four-momentum. Considering that the observer is situated in asymptotically flat spacetime and that the photon is emitted perpendicular to the observation plane, we can safely assume that the initial velocity of the photon, as measured by the observer, is $(\dot{x},\dot{y},\dot{z})=(0,0,1)$. Substituting this into the differential forms of Eqs. (13)-(15), we obtain
$\dot{x}^{\prime}$ $=$ $-\sin\Theta\cos\Phi,$ (21)
$\dot{y}^{\prime}$ $=$ $-\sin\Theta\sin\Phi,$ (22)
$\dot{z}^{\prime}$ $=$ $-\cos\Theta.$ (23)
Subsequently, by inserting $(\dot{x}^{\prime},\dot{y}^{\prime},\dot{z}^{\prime})$ into the differential forms of Eqs. (17)-(19), the components of the photon’s velocity in Boyer-Lindquist coordinates are given by
$\dot{r}$ $=$ $-\frac{r\mathscr{R}\sin\theta\sin\Theta\cos\Psi+\mathscr{R}^{2}\cos\theta\cos\Theta}{\Sigma},$ (24)
$\dot{\theta}$ $=$ $\frac{r\sin\theta\cos\Theta-\mathscr{R}\cos\theta\sin\Theta\cos\Psi}{\Sigma},$ (25)
$\dot{\varphi}$ $=$ $\frac{\sin\Theta\sin\Psi}{\mathscr{R}\sin\theta},$ (26)
@@
Based on Eqs. (8) and (9), it is straightforward to compute $p_{r}$ and $p_{\theta}$ using Eqs. (24) and (25). However, due to the presence of the coupling term $g_{t\varphi}$, the calculation of $p_{t}$ and $p_{\varphi}$ is not as intuitive. Recalling Eq. (7), we have
$\dot{t}=\left(\frac{\Sigma}{\Sigma-2r}\right)\left(-p_{t}-\frac{2ar\sin^{2}\theta}{\Sigma}\dot{\varphi}\right).$ (27)
Substituting $\dot{t}$ into the Lagrangian formula (5) and applying the photon constraint $\mathscr{L}=0$, we obtain
$(-p_{t})^{2}=\left(\frac{\Sigma-2r}{\Sigma\Delta}\right)\left(\Sigma\dot{r}^{2}+\Sigma\Delta\dot{\theta}^{2}\right)+\Delta\dot{\varphi}^{2}\sin^{2}\theta.$ (28)
@@
$p_{\varphi}=\frac{\left(\Sigma\Delta\dot{\varphi}+2arp_{t}\right)\sin^{2}\theta}{\Sigma-2r}.$ (29)

### v2

@@ Abstract
1 Introduction
@@
Since the advent of general relativity, black holes have remained a subject of profound fascination within the astronomical community. In recent years, the study of black holes has evolved from theoretical speculation to direct observation. Notably, the Event Horizon Telescope (EHT) collaboration, utilizing a Very Long Baseline Interferometry (VLBI) array, successfully captured and released the first images of the supermassive black holes at the centers of the M87 galaxy and the Milky Way [1, 2]. These landmark observations not only conclusively verified the existence of black holes but also opened a novel channel for their exploration. Subsequently, a substantial body of research on black hole images [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], black hole-jet system images [18, 19, 20, 21, 22], hot-spot images [23, 24, 25, 26, 27, 28], and polarized images [29, 30, 31, 32, 33, 34, 35, 36, 37, 38], within the frameworks of general relativity and modified gravity theories, has emerged, injecting unprecedented vitality into this field.
@@
Theoretically, a black hole can capture light rays emitted by surrounding sources, resulting in a distinct region of brightness depression on the observation plane from Earth. This region is referred to as the ‘shadow’, first introduced by Falcke et al. [39], and its boundary is the 2D projection of the black hole’s unstable critical photon orbits. Notably, the black hole shadow is considered the ‘fingerprint’ of spacetime, unaffected by the accretion environment, making it a powerful tool for studying black hole physics and gravity theories. Consequently, many pioneering studies have focused on constraining system parameters through the shadow’s profile [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. However, accurately extracting the geometric information of the black hole shadow requires extremely high angular resolution, which is currently beyond reach with the technology of the EHT [4, 52]. As a result, despite the groundbreaking achievements of the EHT, there remains ample room for discussion regarding modifications to gravity theories and the accretion environment.
@@
On the other hand, it is interesting to note that for some supermassive black holes at the centers of galaxies, such as Sagittarius A, the accreting material near the event horizon is optically thin at $230$ GHz [2]. In other words, following the backward ray-tracing approach, the accretion flow within the innermost stable circular orbit (ISCO) can intercept light rays that would otherwise pass through the black hole’s event horizon, allowing the observer to discern a dim region smaller than the traditional shadow. Numerous studies have confirmed that such a dim region can appear when the inner edge of an equatorial, optically and geometrically thin accretion disk coincides with the event horizon, and its boundary is clearly separated from the higher-order bright rings [4, 9, 14, 16, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. More importantly, Chael et al. identified a similar region in general relativistic magnetohydrodynamic (GRMHD) simulations in Kerr spacetime, which they named the ‘inner shadow’ to distinguish it from the traditional shadow [6]. Furthermore, researchers have pointed out that the boundary of the inner shadow is a projection of the black hole’s event horizon, and its geometric informations have the potential to constrain system parameters such as the viewing angle and spin. Building on this, several authors have explored the geometric properties of the inner shadow of Kerr-Melvin black holes illuminated by an equatorial accretion disk across different parameter spaces, finding that the evolution of the inner shadow with varying parameters aligns with the changes in higher-order photon rings. This has led them to propose that the black hole’s inner shadow could serve as a tool for identifying black holes and studying the magnetic field around them [9].
@@
In astrophysics, the angular momentum of accreting matter is not always aligned with the angular momentum of the black hole, which leads to the formation of tilted accretion disk. There is substantial evidence showing that tilted accretion disks are commonly present in X-ray binary systems [64, 65, 66], and the tilted accretion disk-jet system of the supermassive black hole M87 has also been confirmed [67]. Research on the structure evolution and energy spectrum of tilted accretion disks has garnered considerable attention in the scientific community [68, 69, 70, 71, 72]. Inspired by this, Hu and colleagues conducted a simulation of the images of spherically symmetric hairy black holes illuminated by a geometrically thin, tilted accretion disk within the framework of Horndeski theory [14]. They found that the inner shadow of the black hole tends to be obscured by the tilted accretion disk, and it also rotates along with the precession of the disk. Therefore, they argued that, given the difficulties in constraining the accretion environment, using the inner shadow to test gravity theories requires particular caution. However, the scientific community has reached a consensus on the existence of rotation in astrophysical black holes, indicating that Hu et al. did not tell the whole story. In the scenario of a rotating black hole, does a tilted accretion disk obscure the inner shadow? What insights can the inner shadow provide for black hole physics and high-energy physics? These questions will be properly addressed in this paper.
@@
The remainder of this paper is organized as follows. In section 2, we briefly review the Kerr black hole and the canonical equation of photon motion, which form the foundation of the ray-tracing algorithm. Then, in section 3, we provide a detailed introduction to the backward ray-tracing method and the model of the tilted accretion disk. Before the conclusions and discussion in section 5, we present in section 4 the novel inner shadows cast by the Kerr black hole illuminated by a tilted thin accretion disk across different parameter spaces, and carefully elaborate on their astrophysical implications. Throughout this paper, we adopt geometric units, where the black hole mass $M$, the speed of light $c$, and the gravitational constant $G$ are set to unity.
2 Kerr metric and null-like geodesics
@@
Astrophysical black holes are usually described by the Kerr metric, which can be expressed in the Boyer-Lindquist coordinates $x=(t,r,\theta,\varphi)$ with the spacelike signature (,+,+,+)(-,+,+,+)( - , + , + , + ) as [21, 40, 73, 74, 75]
@@
$\textrm{d}s^{2}$ $=$ $-\left(1-\frac{2r}{\Sigma}\right)\textrm{d}t^{2}-\frac{4ar\sin^{2}\theta}{\Sigma}\textrm{d}t\textrm{d}\varphi+\frac{\Sigma}{\Delta}\textrm{d}r^{2}+\Sigma\textrm{d}\theta^{2}$ (2.1)
@@
$\Sigma=r^{2}+a^{2}\cos^{2}\theta,$ (2.2)
$\Delta=r^{2}-2r+a^{2}.$ (2.3)
@@
$r_{\textrm{e}}=1+\sqrt{1-a^{2}}.$ (2.4)
It is important to emphasize that the size of the inner shadow of a black hole is associated with the event horizon [6]. Consequently, it is anticipated that the inner shadow of a Kerr black hole shrinks as the spin parameter increases. Given that there is an upper limit to the spin parameter for a Kerr black hole solution, the size of the inner shadow theoretically has a minimum value. We will delve into this topic in detail in section 4.
@@
$\mathscr{H}=p_{\mu}\dot{x^{\mu}}-\mathscr{L}=\frac{1}{2}g^{\mu\nu}p_{\mu}p_{\nu},$ (2.6)
@@
$p_{t}=-\left(1-\frac{2r}{\Sigma}\right)\dot{t}-\frac{2ar\sin^{2}\theta}{\Sigma}\dot{\varphi},$ (2.7)
$p_{r}=\frac{\Sigma}{\Delta}\dot{r},$ (2.8)
$p_{\theta}=\Sigma\dot{\theta},$ (2.9)
$p_{\varphi}=\sin^{2}\theta\left(r^{2}+a^{2}+\frac{2a^{2}r\sin^{2}\theta}{\Sigma}\right)\dot{\varphi}-\frac{2ar\sin^{2}\theta}{\Sigma}\dot{t}.$ (2.10)
It is necessary to point out that since the Kerr spacetime is of Petrov type D, the metric is independent of the coordinates $t$ and $\varphi$. Consequently, $p_{t}$ and $p_{\varphi}$ are conserved quantities in photon motion, corresponding to the specific energy $E=-p_{t}$ and specific angular momentum $L=p_{\varphi}$, respectively. Hence, we obtain the specific expression of equation (2.6) as
$\mathscr{H}=\frac{1}{2}\left[-\frac{\left(\Delta+2r\right)^{2}-\Delta a^{2}\sin^{2}\theta}{\Delta\Sigma}(-E)^{2}+\frac{\Delta}{\Sigma}p_{r}^{2}+\frac{p_{\theta}^{2}}{\Sigma}+\frac{\Sigma-2r}{\Delta\Sigma\sin^{2}\theta}L^{2}-\frac{4ar}{\Delta\Sigma}(-E)L\right].$ (2.11)
@@
$\dot{x^{\mu}}=\frac{\partial\mathscr{H}}{\partial p_{\mu}},\quad\dot{p_{\mu}}=-\frac{\partial\mathscr{H}}{\partial x^{\mu}},$ (2.12)
@@ 3 Ray-tracing method
Figure 1: Diagram of the ray-tracing method. The local reference frames of the black hole and the observer are denoted by $(x^{\prime},y^{\prime},z^{\prime})$ and $(x,y,z)$, respectively, with the observer’s screen located in the $\overline{xoy}$ plane. $\Theta$ and $\Phi$ represent the observer’s inclination angle and azimuthal angle, respectively. The disk, idealized as a tilted thin accretion disk, is composed of four distinct color-coded quadrants, with its inner boundary connected to the event horizon of the black hole. The tilt angle $\sigma$ is defined as the angle between the black hole’s equatorial plane $\overline{x^{\prime}o^{\prime}y^{\prime}}$ and the tilted accretion disk. Following the positive direction of $\Phi$, the quadrants of the disk are colored red, green, blue, and yellow, respectively. By employing numerical integration algorithms, the trajectories of light rays originating from each pixel $(x,y,0)$ on the observer’s screen can be tracked, with those terminating at the event horizon forming the inner shadow.
As shown in figure 1, the local frames of the observer and the Kerr black hole are represented by quasi-Cartesian coordinate systems $(x,y,z)$ and $(x^{\prime},y^{\prime},z^{\prime})$, respectively. Here, the $z$-axis coincides with $\overline{oo^{\prime}}$ and points towards the point $o^{\prime}$, with an angle $\Theta$ relative to the black hole’s rotation axis ($z^{\prime}$-axis), which defined as the viewing angle. The observer’s screen, depicted as a purple block, lies in $\overline{xoy}$ plane, thus $(x,y,0)$ naturally denotes the initial position of the photon. The solid disk inclined at an angle $\sigma$ to the black hole’s equatorial plane represents a tilted, geometrically thin accretion disk. It is divided into four regions of different colors by quadrants to vividly illustrate the black hole’s inner shadow and the spacetime distortion. It is important to note that for the equatorial accretion disk scenario $(\sigma=0^{\circ})$, the black hole image is independent of the observer’s azimuthal angle. However, this property is expected to not hold when a non-zero $\sigma$ is introduced. Therefore, an additional parameter, the observation azimuth $\Phi$, is included in our model.
Next, we derive the initial position of light rays in the black hole’s local coordinate system. From figure 1, it can be observed that under the condition of asymptotically flatness of spacetime, the observer’s local coordinate system $(x,y,z)$ can be aligned with the black hole’s local coordinate system $(x^{\prime},y^{\prime},z^{\prime})$ through rotation and translation. Thus, the light ray emitted from a pixel $(x,y,z)$ on the observer’s plane can be expressed in the black hole’s local coordinate system as [7, 74, 76]
$x^{\prime}=\mathscr{T}\cos\Phi-x\sin\Phi,$ (3.1)
$y^{\prime}=\mathscr{T}\sin\Phi+x\cos\Phi,$ (3.2)
$z^{\prime}=\left(r_{\textrm{obs}}-z\right)\cos\Theta+y\sin\Theta,$ (3.3)
@@
$\mathscr{T}=\left(\sqrt{r_{\textrm{obs}}^{2}+a^{2}}-z\right)\sin\Theta-y\cos\Theta.$ (3.4)
@@
$r=\sqrt{\frac{\mathscr{D}+\sqrt{\mathscr{D}^{2}+4a^{2}z^{\prime 2}}}{2}},$ (3.5)
$\theta=\arccos\left(\frac{z^{\prime}}{r}\right),$ (3.6)
$\varphi=\textrm{atan}2\left(y^{\prime},x^{\prime}\right),$ (3.7)
@@
$\mathscr{D}=x^{\prime 2}+y^{\prime 2}+z^{\prime 2}-a^{2}.$ (3.8)
To integrate the geodesic equations, in addition to determining the initial position of the photon, it is also necessary to specify its covariant four-momentum. Considering that the observer is situated in asymptotically flat spacetime and that the photon is emitted perpendicular to the observation plane, we can safely assume that the initial velocity of the photon, as measured by the observer, is $(\dot{x},\dot{y},\dot{z})=(0,0,1)$. Substituting this into the differential forms of equations (3.1)-(3.3), we obtain
$\dot{x}^{\prime}=-\sin\Theta\cos\Phi,$ (3.9)
$\dot{y}^{\prime}=-\sin\Theta\sin\Phi,$ (3.10)
$\dot{z}^{\prime}=-\cos\Theta.$ (3.11)
Subsequently, by inserting $(\dot{x}^{\prime},\dot{y}^{\prime},\dot{z}^{\prime})$ into the differential forms of equations (3.5)-(3.7), the components of the photon’s velocity in Boyer-Lindquist coordinates are given by [7, 74, 76]
$\dot{r}=-\frac{r\mathscr{R}\sin\theta\sin\Theta\cos\Psi+\mathscr{R}^{2}\cos\theta\cos\Theta}{\Sigma},$ (3.12)
$\dot{\theta}=\frac{r\sin\theta\cos\Theta-\mathscr{R}\cos\theta\sin\Theta\cos\Psi}{\Sigma},$ (3.13)
$\dot{\varphi}=\frac{\sin\Theta\sin\Psi}{\mathscr{R}\sin\theta},$ (3.14)
@@
Based on equations (2.8) and (2.9), it is straightforward to compute $p_{r}$ and $p_{\theta}$ using equations (3.12) and (3.13). However, due to the presence of the coupling term $g_{t\varphi}$, the calculation of $p_{t}$ and $p_{\varphi}$ is not as intuitive. Recalling equation (2.7), we have
$\dot{t}=\left(\frac{\Sigma}{\Sigma-2r}\right)\left(-p_{t}-\frac{2ar\sin^{2}\theta}{\Sigma}\dot{\varphi}\right).$ (3.15)
Substituting $\dot{t}$ into the Lagrangian formula (2.5) and applying the photon constraint $\mathscr{L}=0$, we obtain
$(-p_{t})^{2}=\left(\frac{\Sigma-2r}{\Sigma\Delta}\right)\left(\Sigma\dot{r}^{2}+\Sigma\Delta\dot{\theta}^{2}\right)+\Delta\dot{\varphi}^{2}\sin^{2}\theta.$ (3.16)
@@
$p_{\varphi}=\frac{\left(\Sigma\Delta\dot{\varphi}+2arp_{t}\right)\sin^{2}\theta}{\Sigma-2r}.$ (3.17)